OpenStudy (anonymous):

Can someone explain to me what goes where? I'm not sure how to solve this..... A metal rod is 25.000cm long at 25.0 degrees Celsius. When heated to 102.0 degrees Celsius, it is 25.054 cm long. What is the coefficient of linear expansion for this metal?

OpenStudy (triciaal):

don't remember exactly but as the temperature goes up the length goes up change in length relates directly to the temperature change in temp = 102 - 25 = 77 degrees change in length = 25.054 - 25.000 = 0.054cm for a change of 77 degrees the length changes 0.054 cm so for 1 degree it changes by 0.054/77 = 7*10^-4cm

OpenStudy (triciaal):

more realistic unit of measure is mm 10 mm = 1 cm expansion is 0.007mm per degree.

OpenStudy (anonymous):

Thank you but I someone explained it to me, like what goes where an I was able to do it myself its 2.805E-5 but thank you :)